1. A stack of nickels and dimes is worth $2.55. Find the number of nickels if the number of dimes is 12 less than twice the number of nickels.
2. Raul is 5 years older than twice Carols age. The sum of their age is 101. How old is Carols?
3. The perimeter of a rectangle is 3 inches less than twice the width. If the perimeter is 18 inches, find the width.

Answer by jojo14344(1513)

Solutions:
1. Let # of nickels = x
So # of dimes = 2x -12, (12 less than twice the # of nickels)
Also when added their total value is $2.55. To show,
0.05(x) + 0.10(2x - 12) = $2.55 -------------- eqn1
continuing,
0.05x+0.2x-1.2=2.55
0.25x = 3.75
x= 15 ------ This is your # of nickels.
In doubt? well just go back eqn1 & substitute x=15, and you get $2.55
2. Let carol's age = x
So, Rauls'age = 2x+5
Adding the their ages will total to 101 years. To show,
x+2x+5=101
3x=101-5
x=32 years ----------- This is Carol's age
So Raul's age= 2(32) +5 = 69 years (an old man ehh?)
Adding the 2 ages, 32+69 =101, cool!
3. Given,
P= 2w-3, (Perimeter is 3 inches less than twice the width right?)
And, P= 18 inches
Equaintg the two conditions,
2w-3=18 ------------ eqn 1
2w=18+3
2w=21
w=10.5 inches (width)
In doubt? Just go back eqn 1 and subsitute w=10.5" and you'll get 18 inches.
Thank you,
Jojo